Testing TI8 started: MSI and MBIH strapped from the access system. The main bends are for the moment...
Download
1 / 3

BTV tested everywhere last week: - PowerPoint PPT Presentation


  • 84 Views
  • Uploaded on

Testing TI8 started: MSI and MBIH strapped from the access system. The main bends are for the moment blocked out. We will have to do the opposite at some point to test the electronic switch (we will run in parallel to CNGS) Asked to have an LHC cycle available all the time after the long MD

loader
I am the owner, or an agent authorized to act on behalf of the owner, of the copyrighted work described.
capcha
Download Presentation

PowerPoint Slideshow about 'BTV tested everywhere last week:' - adrina


An Image/Link below is provided (as is) to download presentation

Download Policy: Content on the Website is provided to you AS IS for your information and personal use and may not be sold / licensed / shared on other websites without getting consent from its author.While downloading, if for some reason you are not able to download a presentation, the publisher may have deleted the file from their server.


- - - - - - - - - - - - - - - - - - - - - - - - - - E N D - - - - - - - - - - - - - - - - - - - - - - - - - -
Presentation Transcript
Btv tested everywhere last week

  • Testing TI8 started: MSI and MBIH strapped from the access system. The main bends are for the moment blocked out. We will have to do the opposite at some point to test the electronic switch (we will run in parallel to CNGS)

  • Asked to have an LHC cycle available all the time after the long MD

  • Warm magnets in point 7 are available with the exception of RD34.LR7. Strapped from the access system this morning. Heat run tomorrow. Try to exercise the warm magnets afterwards waiting for sector 78

Injection Test


Btv tested everywhere last week

  • BTV tested everywhere last week: system. The main bends are for the moment blocked out. We will have to do the opposite at some point to test the electronic switch (we will run in parallel to CNGS)

    • Need to verify that the interlock on the injection screens for B2 is maskable when safe beam flag TRUE (need to have safe beam flag forced to TRUE for point 7 and point 8).

  • Vacuum valves controls, alarms, intlks, logging:

    • LSS6,LSS7 and LSS8, end of TI8 are going to be tested this week.

    • Logging of the vacuum valve status available

    • BIC logging available when the channels are enabled (LSS7 and 8 will be enabled this afternoon) and left enabled.

  • BLM we should really try to get them in BIC and have them running with thresholds

  • Injection inhibit to be tested with LHCb

  • WIC to be included for point 7 and point 8

  • FMCMs?

  • PIC configuration and test for point 7 and 8

Injection Test


Btv tested everywhere last week

  • We should aim at leaving all the channels enabled for the last injection test enabled  Keep the power converters used during the last injection test at injection level

  •  Vacuum valves open in LSS2 and LSS3 during the injection test

  • Status of the Safety Stoppers (strapped from the interlock system so far)

  • Energy from 56-67?

  • Cleaning of Sector 78-81

  • PM generation

  • 5 PM meeting restarting tomorrow

Injection Test